OpenStudy (anonymous):

Give the 3rd and 7th terms using the sequence formula tn = (2n+1)/(3)

jimthompson5910 (jim_thompson5910):

Assuming that the sequence starts at n=1, the 3rd term is \[t_{3}=\frac{2(3)+1}{3}=\frac{6+1}{3}=\frac{7}{3}\] and the 7th term is \[t_{7}=\frac{2(7)+1}{3}=\frac{14+1}{3}=\frac{15}{3}=5\]